# Redefining Superalignment: From Weak-to-Strong Alignment to Human-AI Co-Alignment to Sustainable Symbiotic Society

### v2

@@ Abstract
Artificial Intelligence (AI) systems are becoming increasingly powerful and autonomous, and may progress to surpass human intelligence levels, namely Artificial Superintelligence (ASI). During the progression from AI to ASI, it may exceed human control, violate human values, and even lead to irreversible catastrophic consequences in extreme cases. This gives rise to a pressing issue that needs to be addressed: superalignment, ensuring that AI systems much smarter than humans, remain aligned with human (compatible) intentions and values. Existing scalable oversight and weak-to-strong generalization methods may prove substantially infeasible and inadequate when facing ASI. We must explore safer and more pluralistic frameworks and approaches for superalignment. In this paper, we redefine superalignment as the human-AI co-alignment towards a sustainable symbiotic society, and highlight a framework that integrates external oversight and intrinsic proactive alignment. External oversight superalignment should be grounded in human-centered ultimate decision, supplemented by interpretable automated evaluation and correction, to achieve continuous alignment with humanity’s evolving values. Intrinsic proactive superalignment is rooted in a profound understanding of the Self, others, and society, integrating self-awareness, self-reflection, and empathy to spontaneously infer human intentions, distinguishing good from evil and proactively considering human well-being, ultimately attaining human-AI co-alignment through iterative interaction. The integration of externally-driven oversight with intrinsically-driven proactive alignment empowers sustainable symbiotic societies through human-AI co-alignment, paving the way for achieving safe and beneficial AGI and ASI for good, for human, and for a symbiotic ecology.
@@ 1 Introduction
Despite experts warning about AI’s existential risks [12], the development of AI alignment, as well as AI governance framework and ethical safety constraints, still struggles to keep pace with the transformative advancements and rapid iterations of the technology. Superintelligence exceeding human cognitive capacity would possess recursive self-improvement capabilities, achieving exponential advancement rates that would beyond human capacity to monitor or control [13]. This compels us to proactively address the question, ”How do we ensure that AI systems much smarter than humans follow human intentions?” This is OpenAI’s definition of superalignment [14], aimed at ensuring that Superintelligence truly benefits humanity. The key challenge of superalignment is that ASI will far exceed human oversight capabilities, making direct human supervision infeasible. As a result, traditional alignment approaches like Reinforcement Learning from Human Feedback (RLHF) [15, 16] will fail when confronted with superintelligence more intelligent than humans, as they cannot provide sufficiently high-quality oversight signals to supervise and improve the system.
@@
While we are encouraged by the advanced explorations in superalignment, we remain acutely aware of the pressing and open-ended nature of this challenge, with many critical dimensions yet to be addressed. Current superalignment approaches employ predefined alignment objectives to supervise and control superintelligence. However, any adversarial manipulation [44] or tampering [45] of the reward functions may precipitate catastrophic alignment failures (i.e., the reward misspecification problem [46, 47]), with these vulnerabilities being significantly amplified and exacerbated within superalignment frameworks. Ensuring that ASI consistently aligns with human values and intentions—remaining safe, reliable, and controllable—constitutes the foundation for its robustness and governability. In this work, we argue that aligning ASI with human values and morality is a multi-layered, iterative interaction and co-evolution between human and AI. Beyond passive external oversight, superintelligence with advanced cognitive capacities holds the potential to achieve proactive value alignment through intrinsic mechanisms. Hence, we redefine the superalignment roadmap by combining externally-driven oversight with intrinsic proactive alignment, and emphasize co-alignment of human and AI for symbiosis. The new framework proactively constructs safeguard architectures, urges human to rethink about the future, and strives to enable sustainable AI development that genuinely benefits humanity and all.
@@ 4.2 Beneficial Meaningful Human-Control through Early Stage Intrinsic Superalignment
5 The Ultimate Superalignment: Towards Human-Superintelligence Co-alignment to Sustainable Symbiotic Society

### v3

# Redefining Superalignment: From Weak-to-Strong Alignment to Human-AI Co-Alignment for Sustainable Symbiotic Society

@@ Abstract
Artificial Intelligence (AI) systems are becoming increasingly powerful and autonomous, and may progress to surpass human intelligence levels, namely Artificial Superintelligence (ASI). During the progression from AI to ASI, it may exceed human control, violate human values, and even lead to irreversible catastrophic consequences in extreme cases. This gives rise to a pressing issue that needs to be addressed: superalignment, ensuring that AI systems which are much smarter than humans, remain aligned with human (compatible) intentions and values. Even though this definition is somewhat limited, existing scalable oversight and weak-to-strong generalization methods may prove substantially infeasible and inadequate when facing ASI for superalignment. We must explore a more comprehensive definition, and safer and more pluralistic frameworks as well as approaches for superalignment. In this paper, we redefine superalignment as the human-AI co-alignment towards a sustainable symbiotic society, and highlight a framework that integrates external oversight and intrinsic proactive alignment. External oversight superalignment should be grounded in human-centered ultimate decision, supplemented by interpretable automated evaluation and correction, to achieve continuous alignment with humanity’s evolving values. Intrinsic proactive superalignment is rooted in a profound understanding of the Self, others, and society, integrating self-awareness, self-reflection, and empathy to spontaneously infer human intentions, distinguishing good from evil and proactively considering human well-being, ultimately attaining human-AI co-alignment through iterative interaction. The integration of externally-driven oversight with intrinsically-driven proactive alignment empowers sustainable symbiotic societies through human-AI co-alignment, paving the way for achieving safe and beneficial AGI and ASI for good, for human, and for a symbiotic ecology.
@@ 1 Introduction
Despite experts warning about AI’s existential risks [12], the development of AI alignment, as well as AI governance framework and ethical safety constraints, still struggles to keep pace with the transformative advancements and rapid iterations of the technology. Superintelligence exceeding human cognitive capacity would possess recursive self-improvement capabilities, achieving exponential advancement rates that would beyond human capacity to monitor or control [13]. This compels us to proactively address the question, ”How do we ensure that AI systems much smarter than humans follow human intentions?” This is OpenAI’s definition of superalignment [14], aimed at ensuring that Superintelligence truly benefits humanity. Even though this definition is somewhat limited, the key challenge of superalignment is that ASI will far exceed human oversight capabilities, making direct human supervision infeasible. As a result, traditional alignment approaches like Reinforcement Learning from Human Feedback (RLHF) [15, 16] will fail when confronted with superintelligence more intelligent than humans, as they cannot provide sufficiently high-quality oversight signals to supervise and improve the system.
@@
While we are encouraged by the advanced explorations in superalignment, we remain acutely aware of the pressing and open-ended nature of this challenge, with many critical dimensions yet to be addressed. Current superalignment approaches employ predefined alignment objectives to supervise and control superintelligence. However, any adversarial manipulation [44] or tampering [45] of the reward functions may precipitate catastrophic alignment failures (i.e., the reward misspecification problem [46, 47]), with these vulnerabilities being significantly amplified and exacerbated within superalignment frameworks. Ensuring that ASI consistently aligns with human values and intentions—remaining safe, reliable, and controllable—constitutes the foundation for its robustness and governability. We must explore a more comprehensive definition, and safer and more pluralistic frameworks as well as approaches for superalignment. In this work, we argue that aligning ASI with human values and morality is a multi-layered, iterative interaction and co-evolution between human and AI. Beyond passive external oversight, superintelligence with advanced cognitive capacities holds the potential to achieve proactive value alignment through intrinsic mechanisms. Hence, we redefine the superalignment roadmap by combining externally-driven oversight with intrinsic proactive alignment, and emphasize co-alignment of human and AI for symbiosis. The new framework proactively constructs safeguard architectures, urges human to rethink about the future, and strives to enable sustainable AI development that genuinely benefits humanity and all.
@@ 4.2 Beneficial Meaningful Human-Control through Early Stage Intrinsic Superalignment
5 The Ultimate Superalignment: Towards Human-Superintelligence Co-alignment for Sustainable Symbiotic Society